OpenStudy (anonymous):

find an equation of the line that has the given slope and passes through the given point m=5/3,(3,-1) y=

OpenStudy (anonymous):

start with \[y-y_1=(x-x_1)\] with \[x_1=3,y_1=-1,m=\frac{5}{3}\]

OpenStudy (anonymous):

(y-y1)=m(x-x1)